r=2-2sintheta

theta 0 pi/6 pi/3 pi/2 2pi/3 5pi/6 pi 7pi/6 4pi/3 3pi/2 5pi/3 11pi/6 2pi
r

please show how you worked out each column to find r, i am trying to understand the process
Answer #1

z r=2-2 Sino ufo O then yo 2-asino - 2- 2X0 = 2-0-2 jf @ = 1 then 22-2 Sin (A) = 2-25=2-1=1 IRO=Ę then r7 2-2 sin (I) 2-2x v3if 0= 30 85 2-2 Sin (35) _=-2sin (ET) _ = 2-2 sin( ) =2-2x-1 = 2+2=4 if @ = 50 252-2 šin (55) _ =2-2 Sin (27-5) sin (25-0) =